# Proving membership in LLM pretraining data via data watermarks

### v1

@@ 1 Introduction
Figure 1: An illustration of hypothesis testing for membership inference. The rightholder inserts "MPadd*t6Ex" across their document collection before public release, which was sampled from a distribution of random sequences. The model’s average token loss on all the random sequences forms a null distribution, and the loss on the included watermark is the test statistic. The effectiveness of hypothesis test is determined by the effect size, and variance of the null distribution.
As an example, The New York Times Co. v. Microsoft Corp., S.D.N.Y. 2023222The New York Times Co. v. Microsoft Corp., No. 1:2023cv11195, 17 U.S.C. § 501 Copyright Infringement (S.D.N.Y. filed Dec. 27, 2023). is a recent copyright infringement lawsuit filed in the U.S., where a key piece of evidence is the fact that ChatGPT can reproduce long snippets of historical news articles. [^3] Training on rightholders’ data does not immediately constitute copyright infringement, see Henderson et al. (2023) for an overview of fair use and machine learning. However, whether copyrighted data was included in training will be an important piece of evidence to make a strong case for copyright infringement. In general, using exact reproduction to prove texts were trained on is questionable: certain text may be easy to predict for the model, especially after training on other sources of information on the internet. In pursuing legal action, it would be best to provide statistically sound evidence which indicates that the rightholder’s data was used for training with high probability.
@@ 2 Related work
Tang et al. (2023) use adversarial attacks to create backdoors for verifying dataset membership (coming from a lineage of work on backdoors, see Wallace et al., 2021). However, their methods do not use randomness and remain heuristic. Without randomness, the behavior of a model that has not been backdoored is not easily known. Randomness allows the inserted watermark to be compared against a null distribution of random watermarks; Carlini et al. (2019) study such null distributions for random sequences but in the context of privacy. To the best of our knowledge, our work is the first to combine hypothesis testing and random perturbations to provide principled detection of dataset membership in language models.
@@
Work in membership inference seeks to use the model to infer which data are members of the training data Hu et al. (2022). Our work shares similar goals, but methods in membership inference are primarily motivated by privacy concerns, aiming to extract parts of the training data or sensitive secrets. Our focus is on a narrower case of membership inference, seeking only to detect whether rightholders’ documents were trained on. We assume that the documents could be perturbed beforehand, and our problem admits a statistical solution.
@@
The ability of LLMs to memorize its training data is key to any membership inference. LLMs are known to memorize some of their training data Zhang et al. (2021), and two factors are well-studied relating to an LLMs ability to memorize: the number of times a piece of data is duplicated in the training data (more duplications implies better completion rates; Kandpal et al., 2022), and size of the model (larger models implies better completion; Tirumala et al., 2022b). Our work explores how these key properties subsequently affect the strength of data watermarks.
@@ 3.3 Unicode watermark
We propose a second data watermark that is embedded into the text and imperceptible to humans by using Unicode lookalikes (also called homoglyphs). Unicode attacks are well-studied for a range of text applications and rely on the tokenizer’s sensitivity to Unicode characters to perturb an input sequence Boucher et al. (2021). Further considerations related to watermark stealth are discussed in §7. We curate a conservative list of 28 Unicode lookalike substitutions for the upper and lower case ASCII alphabet. [^7] We use character lookalikes only if they are rendered identically in both Arial and Consolas. These are two default fonts in a popular online browser. The full list of substitutions is given in Appendix C.1. There are two variants of the Unicode watermark: global and word-level, and we instantiate the components of the testing framework below:
@@
For the global Unicode watermark, $\pi$ first generates a random binary vector $v$ of length 28 according to $r$. Each index of $v$ specifies whether the corresponding ASCII character is substituted with its Unicode lookalike everywhere, across all documents in $D$. $\pi$ then returns the document collection where the substitutions specified by $v$ are applied to $D$.
@@
The word-level Unicode watermark replaces each unique word in $D$ with a consistent Unicode lookalike. The perturbation function $\pi$ first creates a random mapping $m$ (seeded by $r$) that maps each unique word $w$ in $D$ (tokenized by whitespace) to a replacement. Each word’s replacement is chosen in the same way as the global Unicode watermark, but we sample a different $v$ for each word and apply it only to that word. Finally, we apply the mapping $m$ to each word in each document $x\in D$ to yield $\pi(D,r)$. Table 1 of the Appendix provides an illustration.
@@ 4 Relating watermark design to strength
In this section, we train many medium-sized language models on watermarked datasets and measure the strengths of the watermarks. We further explore how different properties of the watermark affects the power of the hypothesis test, either by increasing the effect size (i.e., the difference between the test statistic and mean of the null distribution) or decreasing the variance of the null distribution (as illustrated in Figure 1).
@@ 4.1 Experimental setup
We use GPT-NeoX Andonian et al. (2023) to train our language models. The training parameters we use are adapted from Pythia Biderman et al. (2023), which inherits standard practice of training the language model on the shuffled training data for one epoch. This means that each instance of the data watermark is seen only once but its duplications are encountered periodically throughout training. The batch sizes used in this section are small (128 batch size, 512 sequence length) to accommodate the smaller training data.
@@ 4.2 Results
In Figure 3(a), we see that the word-level variant of the Unicode watermark is stronger than the global variant. While the global variant samples one random binary vector of length 28 (the possible Unicode substitutions), the word-level variant samples separate character substitutions for each word. On average, each word has $2.5$ characters that can be substituted for a Unicode lookalike, so the total number of bits is $2.5|V|$, where $|V|$ is the size of the vocabulary (for a collection of 256 documents, $|V|$ is roughly 118,000). Based on the $Z$-scores, this Unicode watermark is nearly equivalent in strength to the 20 length random sequence watermark.
@@ 5 Watermarks under scaling
Both the training datasets and model sizes are very large for popular LLMs. To show that data watermarks can be effective for large-scale models, we fix a watermarked document collection while scaling both the dataset and model size. We focus on the random sequence watermark here, and present similar findings for the word-level Unicode watermark in Appendix C.3.
@@ 5.2 Results
Figure 4(a) shows that watermarks are stronger on larger models, when training on a fixed amount of data. The results here concur with Tirumala et al. (2022a), where they observe that larger models memorize with less epochs. Comparing across 4(b) and (c), the 70M and 410M models have similar null distributions, but the larger model exhibits lower loss on the watermark for the same training setting. The 410M model is qualitatively different than the 70M model under training data scaling, where the test statistic almost does not change at all.
@@ 6.1 Experimental setup
We provide results for the 176B variant of BLOOM Scao et al. (2022) here, and the 7B variant in Appendix D.2. Both models used large batch sizes (512 and 2048, respectively) and were trained on the ROOTS corpus containing 341B tokens for one epoch. The 176B model was trained on an additional repeated 25B tokens.
@@ 6.2 Results
As shown in Figure 5, high-occurrence hashes have lower $p$-values and $Z$-scores, indicating higher watermark strength. However, the correlation between occurrence and strength is imperfect, since occurrences are a weak proxy for the actual number of watermarked documents a hash may have appeared in (many occurrences may be concentrated in one document). [^8] Ideally, we would provide the number of unique documents each hash appeared in, but we encountered limitations of the ROOTS search tool. For instance, we manually confirm that the two SHA-512 hashes with many occurrences but weak $Z$-scores are repeated many times in the same document, so they appeared in fewer distinct training batches than suggested by their number of occurrences. Meanwhile, other hashes in Figure 5 are memorized strongly despite occurring in relatively few documents.
@@ 8 Conclusion
To support a right to opt-out of model training, our work proposes the study of data watermarks. These watermarks provide statistical evidence on whether rightholders’ data has been trained on. By relating aspects of watermark design to the strength of its detection, our insights lay the groundwork for future work on data watermarks. Scaling experiments show that data watermarks are stronger for larger models, and a post-hoc study on naturally occurring SHA hashes confirms that random sequences watermarks could be detected in BLOOM-176B if it occurred more than 90 times in the training data. Together, our results point towards a promising future for data watermarks in real world use.
@@ Appendix A Normality of null distributions
The null distribution is composed of random watermark losses, which are average losses over tokens. The tokens losses may not be independent to each other so the null distributions may not be normal. Normality of the null distribution does not affect the validity of the hypothesis test (see 3.1).
@@ Appendix B Token rarity
A watermark is stronger if it is constructed from rare tokens. To better understand how vocabulary usage affects watermarks, we conduct an oracle study that uses the model’s tokenizer. In Figure 8, we construct our watermarks by sampling random sequences of tokens from different regions of the GPT2Tokenizer which are ordered by frequency (higher rank implies rarer tokens). In particular, instead of always sampling random characters from the first $[0:100]$101010following standard pythonic slicing notation indexes of the GPT2Tokenizer (outlined in Section 3), we randomly sample from the range of $[i:i+100]$, for $i\in\{0,10000,20000,30000,40000,50000\}$. In 8(a), we see that random sequence watermarks composing of rarer tokens have lower Z-scores. Figure 8(b) shows that the test statistic of rarer-token watermarks are lower. We hypothesize that the usage of rarer tokens may induce larger gradient updates during training and exhibit better memorization.
@@ C.1 Unicode lookalikes
C.2 Unicode watermark illustration
An illustration of tokenizations on the word-level Unicode watermark is provided in Table 1.
Seed I have a dream
0 40 423 12466, 108 4320
1 40 289, 16142, 85, 16843 257 288, 260, 16142, 76
2 40 289, 16142, 303 257 288, 260, 16142, 76
Table 1: Tokenizations of the word-level variant of the Unicode watermark by the GPT2Tokenizer. After choosing a seed, each word in the vocabulary is perturbed and we show its corresponding tokenization. Unicode lookalikes can break-up a common word into rare subwords.
C.3 Scaling results
@@ D.2 Results on BLOOM-7B
Figure 9: Test results on naturally occurring SHA and MD5 hashes in BLOOM-7B. Duplication rates are provided by the ROOTS search tool and occurrences may appear in the same document. The dotted line denotes a $Z$-score of $-2$ corresponding to a false detection rate of $\alpha=0.05$.

### v3

@@ 1 Introduction
Figure 1: An illustration of hypothesis testing for membership inference. The rightholder inserts "MPadd*t6Ex" across their document collection before public release, which was sampled from a distribution of random sequences. The model’s average token loss on all the random sequences forms a null distribution, and the loss on the included watermark is the test statistic. The effectiveness of hypothesis test is determined by the effect size and variance of the null distribution.
As an example, The New York Times Co. v. Microsoft Corp., S.D.N.Y. 2023222The New York Times Co. v. Microsoft Corp., No. 1:2023cv11195, 17 U.S.C. § 501 Copyright Infringement (S.D.N.Y. filed Dec. 27, 2023). is a recent copyright infringement lawsuit filed in the U.S., where a key piece of evidence is the fact that ChatGPT can reproduce long snippets of historical news articles. [^3] Training on rightholders’ data does not immediately constitute copyright infringement, see Henderson et al. (2023) for an overview of fair use and machine learning. However, whether copyrighted data was included in training will be an important piece of evidence to make a strong case for copyright infringement. However, most texts used for training cannot be exactly reproduced: while a large journalism organization may have articles heavily duplicated across the internet, most rightholders’ works will not be. For large language models which are trained for one epoch with large batch sizes, a document appearing less than a few times will not be memorized verbatim Duan et al. (2024). In pursuing legal action, it would also be best to provide statistically sound evidence which indicates that a rightholder’s data was used for training with high probability.
@@ 2 Related work
Meeus et al. (2024) is a concurrent work which inserts repeated “copyright traps” in a text document to improve membership inference. Tang et al. (2023) use adversarial attacks to create backdoors for verifying dataset membership. Since these methods do not insert randomness, hypothesis testing cannot be applied. Without randomness, the behavior of a model that has not been backdoored is not easily known. Randomness allows the inserted watermark to be compared against a null distribution of random watermarks; Carlini et al. (2019) study such null distributions for random sequences but in the context of privacy. To the best of our knowledge, our work is the first to combine hypothesis testing and random perturbations to provide principled detection of dataset membership in language models.
@@
Work in membership inference seeks to use the model to infer which data are members of the training data Hu et al. (2022). This literature here has largely been motivated from privacy concerns and aims to extract parts of the training data or sensitive secrets. Recent work finds that membership inference methods perform at the level of random chance on pieces of LLM pretraining data Duan et al. (2024). Crucially, we study a relaxed membership inference setting, as we only seek to detect whether rightholders’ documents were trained on and assume that the documents could be perturbed beforehand. This setting admits a statistical solution and opens up research on membership inference using data watermarks.
@@
The ability of LLMs to memorize its training data is key to any membership inference. LLMs are known to memorize some of their training data Zhang et al. (2021), and two factors are well-studied relating to an LLMs ability to memorize: the number of times a piece of data is duplicated in the training data (more duplications implies better completion rates; Kandpal et al., 2022), and size of the model (larger models implies better completion; Tirumala et al., 2022b). Our work applies these key properties of memorization to detect membership with statistical guarantees, and explores how these properties affect the strength of data watermarks.
@@ 3.3 Unicode watermark
We propose a second data watermark that is embedded into the text and imperceptible to humans by using Unicode lookalikes (also called homoglyphs). Unicode attacks are well-studied for a range of text applications and rely on the tokenizer’s sensitivity to Unicode characters to perturb an input sequence Boucher et al. (2021). Further considerations related to watermark stealth are discussed in §7. We curate a conservative list of 28 Unicode lookalike substitutions for the upper and lower case ASCII alphabet. [^7] We use character lookalikes only if they are rendered identically in both Arial and Consolas. These are two default fonts in a popular online browser. The full list of substitutions is given in Appendix C.1. There are two variants of the Unicode watermark: global and word-level. We instantiate the components of the testing framework below:
@@
For the global Unicode watermark, $\pi$ first generates a random binary vector $v$ of length 28 according to $r$. The random vector’s length of 28 corresponds to the curated list of 28 Unicode lookalike substitutions, where each index of $v$ specifies whether the corresponding ASCII character is substituted with its Unicode lookalike everywhere, across all documents in $D$. $\pi$ then returns the document collection where the substitutions specified by $v$ are applied to $D$.
Seed I have a dream
0 40 423 12466, 108 4320
1 40 289, 16142, 85, 16843 257 288, 260, 16142, 76
2 40 289, 16142, 303 257 288, 260, 16142, 76
Table 1: Tokenizations of the word-level variant of the Unicode watermark by the GPT2Tokenizer. After choosing a seed, each word in the vocabulary is perturbed and we show its corresponding tokenization. Unicode lookalikes can break up a common word into rare subwords.
@@
The word-level Unicode watermark uses $r$ to generate a random binary vector $v_{w}$ for each word $w$ occurring in $D$ (where words in $D$ are delimited by whitespace). Each $v_{w}$ will then be used to substitute all occurrences of $w$ in $D$ with its Unicode lookalike. This is applied in a similar fashion to the global Unicode watermark, but on a word level. An illustration of tokenizations on the word-level Unicode watermark is provided in Table 1. In contrast with the global Unicode watermark, the word-level watermark perturbs each word with a different set of Unicode substitutions, which increases the randomness of the watermark. We investigate this effect along with other differences between global and word-level Unicode substitutions in Section 4.2.
@@ 4 Relating watermark design to strength
In this section, we train many medium-sized language models on watermarked datasets and measure the strength of the watermarks. We further explore how different properties of the watermark affect the power of the hypothesis test, either by increasing the effect size (i.e., the difference between the test statistic and mean of the null distribution) or decreasing the variance of the null distribution (as illustrated in Figure 1).
@@ 4.1 Experimental setup
We use GPT-NeoX Andonian et al. (2023) to train our language models. The training parameters we use are adapted from Pythia Biderman et al. (2023), inheriting standard practice of training the language model on shuffled training data for one epoch. This means that each instance of the data watermark is seen only once but its duplications are encountered periodically throughout training. The batch sizes used in this section are small (128 batch size, 512 sequence length) to accommodate the smaller training data.
@@ 4.2 Results
In Figure 3(a), we see that the word-level variant of the Unicode watermark is stronger than the global variant. While the global variant samples one random binary vector of length 28 (the possible Unicode substitutions), the word-level variant samples separate character substitutions for each word. On average, each word has $2.5$ characters that can be substituted for a Unicode lookalike, so the total number of bits is $2.5|V|$, where $|V|$ is the size of the vocabulary constructed from all words in $D$ (for a collection of 256 documents, $|V|$ is roughly 118,000). Based on the $Z$-scores, this Unicode watermark is nearly equivalent in strength to the 20 length random sequence watermark.
@@ 5 Watermarks under scaling
Both the training datasets and model sizes for popular LLMs are much larger than those we consider in §4. To build intuition for data watermarks when training large-scale models, we fix a watermarked document collection while scaling both the dataset and model size. We focus on the random sequence watermark here, and present similar findings for the word-level Unicode watermark in Appendix C.2.
@@ 5.2 Results
Figure 4(a) shows that watermarks are stronger on larger models, when training on a fixed amount of data. The results here concur with Tirumala et al. (2022a), where they observe that larger models memorize with less epochs. Comparing across 4(b) and (c), the 70M and 410M models have similar null distributions, but the larger model exhibits lower loss on the watermark for the same training setting. The 410M model behaves qualitatively different than the 70M model under training data scaling, where the test statistic nearly does not change at all.
@@ 6.1 Experimental setup
We provide results for the 176B variant of BLOOM Scao et al. (2022) here, and the 7B variant in Appendix D.2. Both models used large batch sizes (512 and 2048, respectively) and were trained on the ROOTS corpus, which contains 341B tokens, for one epoch. The 176B model was trained on an additional repeated 25B tokens.
@@ 6.2 Results
As shown in Figure 5, frequently occurring hashes have lower $p$-values and $Z$-scores, indicating higher watermark strength. However, the correlation between occurrence and strength is imperfect, since occurrences are a weak proxy for the actual number of watermarked documents a hash may have appeared in (many occurrences may be concentrated in one document). [^8] Ideally, we would provide the number of unique documents each hash appeared in, but we encountered limitations of the ROOTS search tool. For instance, we manually confirm that the two SHA-512 hashes with many occurrences but weak $Z$-scores are repeated many times in the same document, so they appeared in fewer distinct training batches than suggested by their number of occurrences. Meanwhile, other hashes in Figure 5 are memorized strongly despite occurring in relatively few documents.
@@ 8 Conclusion
To support a right to opt-out of language model training, our work proposes the use of data watermarks. Rightholders can detect if their data was used for training, by watermarking their data before public release. If the data watermarks are memorized, this serves as statistical evidence on whether rightholders’ data has been trained on. By relating aspects of watermark design to the strength of its detection, our insights lay the groundwork for future work on data watermarks. Scaling experiments show that data watermarks are stronger for larger models, and a post-hoc study on naturally occurring SHA hashes confirms that random sequences watermarks could be detected in BLOOM-176B if it occurred more than 90 times in the training data. Together, our results point towards a promising future for data watermarks in real world use.
@@ 9 Limitations
Acknowledgements
We thank Yiyang Mei for guidance on copyright law and the USC NLP group for their feedback. This work was funded by grants from Open Philanthropy, Cisco Research, and Google Research.
@@ Appendix A Normality of null distributions
The null distribution is composed of random watermark losses, which are average losses over tokens. The tokens losses may not be independent to each other so the null distributions may not be normal. Normality of the null distribution does not affect the validity of the hypothesis test (see §3.1).
@@ Appendix B Token rarity
To better understand how vocabulary usage affects watermarks, we conduct an oracle study that uses the model’s tokenizer. In Figure 8, we construct our watermarks by sampling random sequences of tokens from different regions of the GPT2Tokenizer which are ordered by frequency (higher rank implies rarer tokens). In particular, instead of always sampling random characters from the first $[0:100]$101010following standard pythonic slicing notation indexes of the GPT2Tokenizer (outlined in Section 3), we randomly sample from the range of $[i:i+100]$, for $i\in\{0,10000,20000,30000,40000,50000\}$.
A watermark is stronger if it is constructed from rare tokens. In Figure 8(a), we see that random sequence watermarks composing of rarer tokens have lower Z-scores. Figure 8(b) shows that the test statistic of rarer-token watermarks are lower. We hypothesize that the usage of rarer tokens may induce larger gradient updates during training and exhibit better memorization.
@@ C.1 Unicode lookalikes
C.2 Scaling results
@@ D.2 Results on BLOOM-7B
Figure 9: Test results on naturally occurring SHA and MD5 hashes in BLOOM-7B. Duplication rates are provided by the ROOTS search tool and occurrences may appear in the same document. The dotted line denotes a $Z$-score of $-2$ corresponding to a false detection rate of $\alpha=0.05$. Since the model is relatively small to the dataset, more duplications are needed for detection.